Displaying reports 941-960 of 1092.Go to page Start 44 45 46 47 48 49 50 51 52 End
Reports until 16:57, Friday 09 February 2024
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 16:57, Friday 09 February 2024 - last comment - 17:00, Friday 09 February 2024(11345)
Noliac NAC2125-H08 (Ring Piezo) Testing

Alex and Daniel

We checked a Noliac NAC2125-H08 for shorts and found none. We measured the capacitance to be 2.4641 uF. The expected capacitance is 2.4 uF. We also measured a resistance of 3.5082 kOhm. When we pressed on the piezo, we noticed a voltage. We flagged one of the wires for polairty. See attached photos. We also noticed that the nylon tipped set screws that hold the piezo in place provide a voltage.

Images attached to this report
Comments related to this report
Daniel Grass - 14:34, Thursday 26 October 2023 (11354)

I looked for the dots on the Piezo we tested, and they are on the wire we DID NOT FLAG. I think we should flip which wire is flagged and then use that convention going forward.

Daniel Grass - 17:00, Friday 09 February 2024 (11355)

I flagged the lead of the piezo with the black dot, which is explicitly stated by Noliac as the positive electrode. When I press on the piezo, there is a positive voltage when measuring off this positive lead.

Images attached to this comment
GQuEST Output Filter Cavities
Daniel Grass - posted 17:27, Monday 30 October 2023 (11360)
Newport Mirrors in Bowtie Subassemblies

I put one Newport Dielectric Mirror, Laser Line, 25.4 mm, λ/10, 1520-1580 nm (10D20DM.8) each in the crusher and piezo bowtie subassemblies. The piezo subassembly also has a Clean Room Viton Fluoroelastomer O-Rings, Chemical-Resistant, 1/16 Fractional Width, Dash Number 020, 1309N22, from McMaster. After compressing the piezo and the o-ring by threading in the SM1 ring, I backed off the screws that hold the piezo.

The subassemblies themselves are mounted to test. I set up a Michelson Interferometer to test the piezo. The piezo leads are connected to a BNC, with the positive lead in the center of the BNC and the negative lead to ground. A Thorlabs MDT693B - 3-Channel, Open-Loop Piezo Controller is also ready to be used. The output port of the Michelson is a 1550 nm photodiode.

The crusher is mounted on a bowtie cavity and the beam profiler is ready to be used there.

Right now, I am waititng on a Thorlabs order I placed today to get fiber outputs mounted into SM1 rings.

SNSPD General (GQuEST)
Alex Ramirez - posted 15:42, Friday 27 October 2023 (11358)
Thermal Source Assembly for temperature dependence testing

I have dissasembled the axetris - EMIRS50 AT06V BR25M thermal source such that the reflective can was desoldered. Next, to properly mount the thermal source into the previously described lens tube assembly (11334) the thermal source was soldered onto a brass washer using thermal paste and the PCB oven. The end results are shown bellow in attachments 1-3. The parts for the assembly will be here on monday to be machined and assembled.

The source was placed as closely to the center of the washer as possible, but a custom washer may need to be designed such that minimum alignment is needed when assembling the fiber coulped device. 

The SNSPD will undergo a PCR test and be placed in the fridge monday as well. 

Images attached to this report
GQuEST Output Filter Cavities (GQuEST)
Daniel Grass - posted 15:40, Friday 27 October 2023 (11359)
Newport AJS254-0.5K-NL Adjustment Screw Dip Dissolution in Isopropyl Alcohol and Sonication

While cleaning the Newport AJS254-0.5K-NL Adjustment Screw, I sonicated three Newport AJS254-0.5K-NL Adjustment Screws in isopropyl alcohol at 30°C for 20 minutes. When I came back, I discovered the tip had dissolved. I thought the tip was stainless steal like the body, but it might not be. I reached out to Newport for clarification. If the crusher does work (we at least one screw on hand), we will need 3 more replacements.

GQuEST Output Filter Cavities
Torrey Cullen - posted 16:35, Wednesday 25 October 2023 (11353)
Readout Filter cavity is ready for science

[Daniel Torrey]

One of the low power paths in the power distribution center (See "C:\Users\gques\Nextcloud\GQuEST\Layout_Mockups\B102_starting_setup_pwr_dist.svg" for updated layout) is aligned with 84% efficiency through the fiber. We have light available for readout filter cavity science.

Lasers General
Torrey Cullen - posted 16:15, Wednesday 25 October 2023 - last comment - 10:13, Tuesday 31 October 2023(11352)
SHG profile

[Daniel Torrey]

Profiled the output of the SHG, data for which can be found at Nextcloud/GQuEST/B102/SHG/SHG Beam Profile/ .See attached. Note the output for the SHG fiber has a collimator, very easy to change the beam profile.

Images attached to this report
Comments related to this report
Torrey Cullen - 10:13, Tuesday 31 October 2023 (11362)

Note that this profile is outdated, I adjusted the output collimator so that the beam would fit through the small aperture of the EOM. Updated profile should be in the next cloud B102 directory.

Lasers Amplifiers
Torrey Cullen - posted 16:29, Tuesday 24 October 2023 - last comment - 14:02, Wednesday 25 October 2023(11348)
SHG efficiency discrepancy

Set current to 2.3 amps. This corresponds to 3.96 W out of the amplifier. The beam goes through a PBS, 50:50, 90:10, and PBS. The waveplates before the PBS can be optimized to give <1% loss in the S pol. Assuming minimal loss at the 50:50 and 90:10 this gives 1.75 W into the fiber. With 85% efficiency through the fiber my best guess is we have ~1.48W going into the SHG. The readout on the power meter shows 38 mW. This can be adjusted by the ratios in responsivities of the power meter at 775 and 830 so 38mW * 207.1/146.77 = 54 mW of 775 nm light. If we believe the covesion data sheet it predicts ~350 mW for similiar inputs. Could our SHG be that much more inefficient? Based on the power we ARE getting out of the SHG, according to the data sheet, you would think were feeding in approximately 500 mW, which I don't think is the case. Don't want to stick the power meter in there though.

2.3 A on the amplifier was the previously calculated acceptable operating point to not damage the faraday isolator at the amplifier output. I'm assuming 50 mW of 775 nm light is not sufficient for all of gquest, although I don't have a good sense for this number.

Comments related to this report
Torrey Cullen - 14:02, Wednesday 25 October 2023 (11351)

Lee suggested to check we are at the correct temperature. According to this test the peak temp is 50.423 C, whereas the manual suggests 50.461. There is a slight uncertainty since we don't have a way to perfectly sync the time scales of the moku and temperature controller logger. Even still though the efficiency as a function of temperature is fairly flat near the peak. I don't think this explains the big discepancy in efficiency of the SHG.

Images attached to this comment
GQuEST General (GQuEST)
Daniel Grass - posted 15:59, Tuesday 24 October 2023 (11347)
Noliac Info on Wiki

https://wiki.mccullerlab.com/Main/NoliacPiezoInfo

Lasers General
Torrey Cullen - posted 12:28, Monday 23 October 2023 - last comment - 15:05, Monday 23 October 2023(11340)
SHG first light

SHG and temperature controller turned on and we are seeing 775 nm light (probably). Couple of things:

1) The thorlabs power meter doesn't have a 775 nm light calibration, if we want to detect 775 we may want to find another power meter. The closest it has is 830 nm. 

2) SHG temp controller takes ~350 seconds to stablize to within .005 of a degree. Set it to 50.461 C for maximum 775 effeciency.

3)The SHG cavity it wildly inefficient at low input powers, it produces 38.5 uW of 775 with an input of 28 mW (.13% effeciency, also for the record 4 uW measured at 1550 as well). I ramped the current on the amplifier to 1A. This corresponds to an input power of 93 mW which yields an output of 150uW (.16% effeciency). The SHG has been assigned the serial number S2300012, see that link for specs and manual for this which has an efficiency curve as a function of input power. Similarly for the temperature controller S2300013.

We also need a more permanent way to mount it to a breadboard if anyone wants to 3D print something for it.

Comments related to this report
Torrey Cullen - 15:05, Monday 23 October 2023 (11341)

Daniel, Torrey, and Sander

To measure the power out of the SHG, we are using a Thorlabs S132C without the ND filter on. We measured the output of the SHG before and after reflection at a 45° AOI off a Thorlabs BB1-E03 mirror. The BB1-EO3 reflects p-polarized 775 nm light at 99.2 and 1550 nm light at 7.0%. There was no noticible change in the power before and after reflection, indicating very little 1550 nm was transmitted out of the SHG. To measure the 775 nm power from the SHG, we set the S132C to 700 nm and 830 nm. Using the given responsivity (mA/W) specs as a function of wavelength, we can convert the given power at these selected wavelenghts to 775 nm power. For a setting of 830 nm, multiply the given power by 1.41 to get the 775 nm power. For a setting of 700nm, multiply the given power by 0.54 to get the 775 nm power. Doing this gives 1.59 mW and 1.61 mW, respectivly, for infered 775 nm power. We were expecting around 5 mW for a 200 mW input based on the SHG specs, but the graph is difficult to read. Further analysis is needed.

Images attached to this comment
GQuEST General (GQuEST)
Daniel Grass - posted 22:49, Saturday 21 October 2023 (11339)
Heater Requirements for the GQuEST Cryostat at Equilibrium

Since the pulse tube cannot be easily turned on and off, I added a few lines of code to determine how much power the heaters on the coldplate must emit at equilibirum to maintain a temperature of 123 K. For the Cryomech PT30RM, it is 11 W, while for the Sumitomo CH110, it is 19 W. There is a small risk that this value is an order of magnitude too low for the CH110 since it has a lot more cooling power at higher temperatures. I belive my modeling is correct, howerver. For the PT30RM, I belive 11 W is close to the actual value.

 

I think this implies we should have a high power heater near the coldhead's linkage to provide a rough constant temperature and a much smaller power heater on the beamsplitter mount to provide fine PID control.

 

See attached file

Non-image files attached to this report
GQuEST General
Daniel Grass - posted 12:43, Friday 20 October 2023 (11337)
Using a Thermal Reservoir to Cool the Beam Splitter without Vibrations

While considering vibrations of the coldhead and what to do if they're problematic for the interferometer, an idea occured to me to cool the conductive bar as a reservoir, turn off the cold head, and then the beam splitter will slowly cool down from 123 K to 122 K and then heat up to 124 K. In order to tune this temperature, we would need to in situ remove ~95% of the conductivity from the bar to the cold breadboard. This remaining conduction would have to be pretty accurate as well. I consider a 10 kg copper reservoir and the two "best" coldheads, the Sumitomo CH110 which has a lot of cooling power, and the Cryomech PT30RM which has a fair amount of cooling power but should have minimal vibrations. In summary, this is a vialbe option if we can do two challeneing things: change the conduction between the bar and the cold breadboard in situ and turn the coldhead on and off a few times per day.

 

In addition, I looked into two other tricks to help us: increasing the mass of the cold breadboard and thermally linking the cold breadboard and the inner shield to provide additional time. The former does help the duty cycle and increases the warmup time while the latter has a worse duty cycle with minimal benefit over just cooling more mass. 

 

Modeling data below:

 

Hot inner shield, 4 kg cold breadboard:

CH110:

6:09:57 to cool everything initially

Get 7:33:57 before the breadboard hits 124 K, takes 0:04:39 to reset breadboard, 34:58 to reset bar to within 1 K of equilibrium (duty cycle = 0.94)

 

PT30RM:

13:54:56 to cool everything initially

Get 6:08:26 before the breadboard hits 124 K, takes 0:10:43 to reset breadboard, 4:03:00 to reset bar to within 1 K of equilibrium (duty cycle = 0.6)

 

Hot inner shield, 20 kg cold breadboard:

CH110:

26:09:36 to cool everything initially

Get 15:42:36 before the breadboard hits 124 K, takes 0:24:30 to reset breadboard, 0:45:37 to reset bar to within 1 K of equilibrium (duty cycle = 0.95)

 

PT30RM:

44:28:12 to cool everything initially

Get 11:45:47 before the breadboard hits 124 K, takes 1:14:48 to reset breadboard, 4:56:26 to reset bar to within 1 K of equilibrium (duty cycle = 0.70)

 

Cold inner shield, 4 kg cold breadboard:

CH110:

23:49:17 to cool everything initially

Get 17:57:21 before the breadboard hits 124 K, takes 3:55:45 to reset the breadboard, 12:22:07 to reset the inner shield to within 1 K of equilibrium (duty cycle = 0.6)

 

PT30RM:

43:15:39 to cool everything initially

Get 16:09:28 before the breadboard hits 124 K, takes 9:00:23 to reset breadboard temp to 123 K, 15:02:13 to reset the inner shield to within 1 K of equilibrium (duty cycle = 0.5)

Non-image files attached to this report
Lasers Amplifiers
Torrey Cullen - posted 15:40, Thursday 19 October 2023 - last comment - 17:13, Friday 20 October 2023(11336)
Input coupler test

[Alex Torrey]

We did a quick test of the input couplers we have on hand. Initially couldn't get more than 1% through the fiber. Plopped a f=1m lens about a meter away from the input coupler and now get ~80%. Will com eup with a proper MM solution for this tomorrow. 

 

Images attached to this report
Comments related to this report
Torrey Cullen - 17:13, Friday 20 October 2023 (11338)

The input coupler was not in its final position so we moved it and recovered ~80% transmission through the fiber. We discovered that how the input couplers are mounted currently is not fully recoverable. We will try turning on the SHG on monday and when the proper mounts adapters and mounts arrive (tuesday) we will put the input couplers in the new mounts and recover the fiber alignment.

SNSPD General (GQuEST)
Alex Ramirez - posted 14:43, Wednesday 18 October 2023 (11334)
SNSPD Thermal Response Testing

Boris and I have been working towards getting the SNSPD into the cryo chamber. There have been some hiccups as the chambers in Downs 123 have gone down recently, but we are looking toward taking an initial PCR curve next week (10/23/23) for the newly assembled SNSPD. See images 1-4 for the constructed SNSPD, wirebonded at JPL.

For the initial PCR curve on the SNSPD (to ensure it is working properly) we will  use the top seen in image 3, which will allow us to place a fiber directly above the SNSPD for testing. Next, once we are happy with the results, we will go ahead and replace the top with a dark box enclosure to test the SNSPD performance on dark counts while completelty enclosed in the cryo-chamber.

Finally, we will be testing the SNSPD's thermal response to radiation using a thermal source (axetris - EMIRS50 AT06V BR25M, attachment 7) and the setup seen in image 5. This setup will consist of the voltage controlled thermal source being placed approximately 1 mm away from the fiber tip allowing for complete exposure. The remaining fiber will be encapsulated in a brass tubing leading to the SNSPD. The setup utilizes thorlabs SM05 tubes and brass optic spacers to allow the thermal source to be held closely to the fiber. See image 6 for a real image of this thermal source.

We will be ordering these parts soon and assembling in the coming weeks. 

Images attached to this report
Non-image files attached to this report
GQuEST Output Filter Cavities
Daniel Grass - posted 17:14, Tuesday 17 October 2023 (11333)
Crusher Data Analysis

Torrey and Daniel

We analyzed the data that I took from crushing the mirror. The results are a little weird; crushing the mirror didn't move the horizontal waist location but did change its size. The horizontal waist location got moved forward which is the incorrect direction. I pushed on the mirror as hard as I could and didn't break it, so I think we should try a coated mirror. 

 

Files attached

Non-image files attached to this report
Lasers Amplifiers (GQuEST, Layouts)
Daniel Grass - posted 15:41, Tuesday 17 October 2023 - last comment - 12:01, Tuesday 24 October 2023(11330)
Setting up the High Power Distribution Center

[Torrey Daniel]

Began initial set up of the power distribution center. We have all the parts according to the diagram. Things are roughly aligned up until the first 50:50 BS. Will do alignment into input fiber couplers tomorrow. Torrey will attach photos and diagrams in a comment on this post.

We also put all of the laser initialization on one breadboard. This includes the seeder, polarization paddles, fiber PBS, and amplifier. The amplifier output is mounted on the power distribution center.

Comments related to this report
Torrey Cullen - 15:55, Tuesday 17 October 2023 (11331)
Images attached to this comment
Torrey Cullen - 10:14, Thursday 19 October 2023 (11335)

Update to power distrubtion center. We have things roughly aligned. The four large 1/4-20s at the end of each path represent an equal distance from the first 50:50 BS. We need more fiber collimators/input couplers however. There are 3 of these which seem less than ideal for input alignment as they have no way to control the z-axis. I suggest we get more of Lee's prefered collimator. At least three but more would be better for the future. Also going to buy better ways to mount the F230APC-1550's.

Non-image files attached to this comment
Torrey Cullen - 15:28, Monday 23 October 2023 (11342)

The beam splitters were mounted a 1/4 inch too high. Remounted and did course alignment. We also have a proper way to mount the fiber input couplers (they were slipping). Replaced those and put them in their approximate location. We need to make a proper mode matching solution. I don't know the exact beam size these collimators want though.

Torrey Cullen - 12:01, Tuesday 24 October 2023 (11346)

RE: need a proper MM solution for this path.

I double checked what the collimator beam looks like coming out from old data, the spec sheet quotes 530 um, our measured data was ~460 um. A 1 m focal length lens roughly 1 meter away from the collimator inputs provides a 94% match. Keeping it like this. Attached the jam file and a picture in case you don't have jammt.

Images attached to this comment
Non-image files attached to this comment
Displaying reports 941-960 of 1092.Go to page Start 44 45 46 47 48 49 50 51 52 End